Home CMS Production Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  Belle testbeam at SPS in June 2015, Page 6 of 6  Not logged in ELOG logo
ID Date Author Project Subject Run Number Events XYPosition Cooling Data Text Attachments
  7   Thu Jun 4 18:44:50 2015 Hao YinBelle II FADC (H6A)Hardware002500002NoGoodroom temp, FIRRun001, pedestal run 003  19x 
  6   Thu Jun 4 15:12:41 2015 Hao YinBelle II FADC (H6A)Hardware001500001NoGoodPedestalRun001, FIRRun001, room temp
telescope n-side switched.
In order to correct the position change
  5   Thu Jun 4 14:16:20 2015 Hao YinBelle II FADC (H6A)Pedestal00110000  GoodFirRun001, room temp
telescope n-side switched.
  4   Thu Jun 4 14:15:50 2015 Hao YinBelle II FADC (H6A)FIR001   GoodFIR filer and ADC delay with all sensor
and telescope.
telescope n-side switched.
  3   Wed Jun 3 17:23:07 2015 Thomas BergauerBelle II FADC (H6A)XY stage positions    Unknownx-Achse: positiv Richtung Saleve (counting
  2   Wed Jun 3 10:48:42 2015 Benedikt WürknercommonNetworking    UnknownWLAN Router:


password: 20supersecret14
  1   Tue May 12 10:27:39 2015 Benedikt WürknerSVD3 (H6B)Setup information    UnknownInformation concerning the SVD3 setup during
a beam test

In the beam area setup the whole
ELOG V3.1.4-395e101